OpenStudy (abbycross167):

can someone please help me with this math question? please I'll fan and give a medal...

OpenStudy (abbycross167):

show the steps to graph this equation -- x + y <= -3

OpenStudy (abbycross167):

@jim_thompson5910 would you mind helping me again?

jimthompson5910 (jim_thompson5910):

how far did you get with this one?

OpenStudy (abbycross167):

no where... I'm not sure how to do it since both the x and y have no numbers

jimthompson5910 (jim_thompson5910):

first, let's replace the inequality sign with an equal sign so `x+y <= -3` will turn into `x+y = -3`

jimthompson5910 (jim_thompson5910):

now let's find the y intercept to do this, replace every copy of 'x' with 0. Then isolate y

jimthompson5910 (jim_thompson5910):

what do you get when you do this?

OpenStudy (abbycross167):

so it would look like this right? -3+y?

jimthompson5910 (jim_thompson5910):

x+y = -3 0+y = -3 ... replace every 'x' with 0 y = -3 .... simplify

jimthompson5910 (jim_thompson5910):

So when x = 0 this makes y = -3 They pair up to form the ordered pair (0,-3) which is the y intercept

OpenStudy (abbycross167):

okay, would x be the same?

jimthompson5910 (jim_thompson5910):

x-intercept x+y = -3 x+0 = -3 .... replace every 'y' with 0 x = -3 .... simplify yes that is correct. The x-intercept would be (-3,0)

jimthompson5910 (jim_thompson5910):

plot the two points, then draw a solid line through them then determine where the shaded region should be based on a test point. (0,0) is often a good test point to try out

OpenStudy (abbycross167):

Thank you so much!! @jim_thompson5910 you are very very helpful!!!

jimthompson5910 (jim_thompson5910):

I'm glad I could help